OpenStudy (anonymous):

Let (x y^2) - (x^3 y)=6 and m = ((3x^2 y)-(y^2)) / ((2xy)-(x^3)). Find all points (x,y) on the curve where the slope is of the tangent line is horizontal, then vertical?

OpenStudy (anonymous):

is that m = dy/dx ?

OpenStudy (anonymous):

they already gave you the derivative or is that your calculated derivative?

OpenStudy (anonymous):

this is the problem I was given. I'm beyond confused

OpenStudy (anonymous):

but is that m = dy/dx?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

ok...let's compute where the tangent is horizontal..

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

what would the slope be if the line is horizontal?

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

what's the slope of a horizontal line?

OpenStudy (anonymous):

yes... zero...

OpenStudy (anonymous):

okay so (0, something

OpenStudy (anonymous):

right... just set m=0 and you'll be solving this equation: \[\huge 0=\frac{3x^2y-y^2}{2xy-x^3} \]

OpenStudy (anonymous):

so essentially only the numerator will make this equation true so you're actually solving this: \[\huge 0=3x^2y-y^2 \]

OpenStudy (anonymous):

Okay....what exactly am I solving for? I'm sorry I'm having such a hard time. I've been working on 204 calculus/pre-calculus review problems. My brain is fried and I don't remember quite a few things.

OpenStudy (anonymous):